OpenStudy (anonymous):

The total cost for 8 bracelet, including shipping was $54. The shipping charge was $6. Determine the cost for each bracelet.

OpenStudy (jim766):

cost of braclet = B 8 * cost of braclet + $6 shipping = 54 8b + 6 = 54, now solve for b

OpenStudy (jim766):

subtract 6 on both sides, then divide by 8 on both sides

OpenStudy (anonymous):

I still don't understand the question

OpenStudy (jim766):

first you need to translate the words into math 8 * cost of 1braclet + $6 shipping = 54 we don't know what the price of a braclett is so we call it B 8 * B + 6 = 54 -6 -6 8B = 48 now divide both sides by 8 8B = 48 -- --- 8 8 B = 48/8 what is 48 divided by 8?

OpenStudy (anonymous):

48 /8=6

OpenStudy (jim766):

yes, B = 6, or $6 the cost of 1 braclet is $6